# A Survey on Causal Discovery: Theory and Practice

### v1

@@ Abstract
Understanding the laws that govern a phenomenon is the core of scientific progress. This is especially true when the goal is to model the interplay between different aspects in a causal fashion. Indeed, causal inference itself is specifically designed to quantify the underlying relationships that connect a cause to its effect. Causal discovery is a branch of the broader field of causality in which causal graphs is recovered from data (whenever possible), enabling the identification and estimation of causal effects. In this paper, we explore recent advancements in a unified manner, provide a consistent overview of existing algorithms developed under different settings, report useful tools and data, present real-world applications to understand why and how these methods can be fruitfully exploited.
@@ 1.1 A General Overview
One of the mantras that is repeated in every statistical course is that correlation does not imply causation. This is also observed in several disciplines, such as economics [38], biology [68], computer science [33, 71] and philosophy [27]. Following [32], the main goal of a research study is often to assess the effect, if any, of an action on some outcome and not measuring a mere correlation. For example, this is true when it comes to decision making, since deciding which intervention must be taken is not straightforward and must be addressed properly to avoid any potential side effects. In order to identify and quantify a causal effect, the set of tools provided by causal discovery must be used accordingly. Here, the final goal is to decompose the total effect of an action into causal and non-causal effect, removing the bias that is introduced during the estimation process.
Causal inference itself relies heavily on a formal description on the interactions between the observed variables, i.e. a casual graph. Such graphical representation is naïve in its concept, yet so effective when it comes to explainability. Following [99], it boils down to connect a cause to an effect (outcome) by drawing arrows from the former to the latter, to obtain a qualitative description of the system under study. This is in stark contrast with black-box techniques, where predictions about an outcome are made with a pure data-driven approach. Indeed, these methods fall short both in terms of explainability and decision making, as stated in [8, 28, 32]. Therefore, when causality is empowered through the instrument of graphical models, it is possible to overcome the current limitations of machine learning and deep learning tools, enabling the researcher to reach a higher level of understanding.
@@
This survey is structured as follows. In Section 1, the reader is provided a general introduction to the causal discovery problem, along with an overview of previous works on the same topic. Section 2 is devoted to provide concepts, definitions and problems that are common across different approaches presented in the following pages. Section 3 explores the first set of algorithms in the observational setting, while Section 4 relaxes the acyclicity assumption. In Section 5, the scope is extended to cover the experimental scenario, where multiple interactions with the system of interest are taken into account. Section 6 and 7 report respectively on evaluation techniques and on practical applications of the discussed methodologies. Finally, Section 8 draws conclusions about the current landscape of causal discovery.
@@ 1.2 Related Works
To the best of our knowledge, six different surveys on causal discovery were published from 2019 to 2022. In particular, [67] acted as a meta-survey by checking the contents covered by the others against five topics, namely: theory, data, software, metrics and examples. A modified version of this checklist can be found in Table 1, which was adapted for a direct comparison with the structure of our survey.
@@ A Survey of Learning Causality with Data: Problems and Methods [29]
While every contribution provided adequate background knowledge and theoretical definitions involving the fundamental aspects of causal discovery, only few examples [27, 63, 66] reported evaluation data sets or metrics, and just two of them listed both [63, 67]. The landscape is even more fragmented when observed from a practical point of view: only two contributions [66, 67] presented and discussed the availability of software tools to perform the described procedures, thus hindering the applicability of causal discovery to researchers approaching for the first time to this topic.
In particular, the work from [29] provides some insights on the discovery procedure from a deep-learning point of view. Authors in [63] tackled the problem of recovering the causal graph from time series data sets, while [27] restricted its attention to the most famous techniques. Moreover, [66] presented a general survey on the topic without a proper interventional section, as for [86] in the latent case. Finally, [67] covers causal inference and causal discovery from a high-level perspective, which is opposed to our in-depth approach focused on structural learning only.
This survey is designed in the light of the above considerations and aims to guide the inexpert reader through the forest of causal graphs to avoid common pitfalls when comparing and assessing the quality of results obtained by different causal discovery algorithms. It is worthwhile to mention that this survey is different from those published from 2019 to 2022 with respect to both theory and practice. Indeed, existing surveys introduce theoretical aspects of causal discovery while only few of them go into additional details. Another lack of existing surveys in term of theory is that very few of them discuss the difference between observational and interventional data. This survey has also many differences in terms of practice; i) Only a subset of existing surveys report on evaluation datasets and metrics, ii) no survey discusses how to tune strategies for choosing values of algorithm’s hyperparameters, iii) only one survey reports on software packages, and iv) few surveys discuss practical applications of causal discovery methods.
@@ 2 Definitions and Notation
This section gives the main definitions, concepts and assumptions on causality, together with the associated notation. In particular, we give the definition of causal model and the definition of causal discovery problem.
@@ 2.1 Notation
[…]
@@
In particular, a directed edge (X, Y) is graphically represented by an arrow as $X\rightarrow Y$, and induces a set of relationships between the vertices of the graph $G$. Given a vertex $X$, we denote its parents, i.e., the set of vertices that have an arrow into $X$, by $Pa(X)$, while we denote its children, i.e., the set of vertices that have an arrow out of $X$, by $Ch(X)$. Recursively, any parent and parent of a parent (child and child of a child) of $X$ is an ancestor $An(X)$ (descendant $De(X)$) of $X$.
The vertices connected to $X$ are said to be adjacent to $X$ and denoted as $Adj(X)$, while the vertices connected with an undirected edge to $X$ are the neighbors $Ne(X)$. These two sets of vertices are identical in undirected graphs, but may be different in graphs with other mixed orientations.
@@
A path $\pi=(X-\cdots-Y)$ is a tuple of non repeating vertices, where each vertex is connected to the next in the sequence with an edge.
@@ 2.2 Causal Model
A causal graph $G$ [8] is a graphical description of a system in terms of cause-effect relationships, i.e. the causal mechanism.
@@
This definition is formally enforced by the causal edge assumption [28], where:
@@
The value assigned to each variable $X$ is completely determined by the function $f$ given its parents:
$X_{i}:=f(Pa(X_{i}))\qquad\forall X_{i}\in\textbf{V}.$ (2.1)
As natural consequence of such definitions, we can define models that entail both the structural representation and the set of functions that regulate the underlying causal mechanism.
@@
  • U is a set of exogenous variables, i.e. unobservable variables, where $\textbf{V}\cap\textbf{U}=\varnothing$,
@@
$M$ $=(\textbf{V},\textbf{U},\textbf{F},P)$ V =$=\{X,Y,Z\}$ U =$=\{U_{XY},U_{Z}\}$ F =$=\begin{cases}f_{X}:X:=2U_{XY},\\ f_{Y}:Y:=X+U_{XY},\\ f_{Z}:Z:=3Y+U_{Z}\end{cases}$ $P$ =$=\begin{cases}U_{XY}\sim\mathcal{N}(0,1),\\ U_{Z}\sim\mathcal{N}(0,1)\end{cases}$
Figure 2.1: The causal graph $G$ (a) of the related SCM $M$ (b). In (a) $X$ is a direct cause of $Y$ and an indirect cause of $Z$, while $Y$ is an effect, a direct effect, of $X$. An example of associated SCM is reported in (b), where the functional set F follows the causal edge assumption.
@@ 2.3 The Causal Discovery Problem
Formally, let G be the set of graphs defined over the variables V of a data set D and $^{*}\in\textbf{G}$ G be the true but unknown graph from which D has been generated.
@@
A causal discovery algorithm is said to solve the causal discovery problem if and only if it converges to the true graph $G^{*}$ in the limit of the sample size.
@@
A causal discovery algorithm is sound if it is able to solve the causal discovery problem, while it is complete if it outputs the most informative causal graph $G$ that can be recovered from the input data set D, without making further assumptions.
@@
In the following pages we will see that some algorithms are able to identify the causal graph up-to its equivalence class, meaning that setting the direction of any of the remaining undirected edges would not induce a different probability distribution, i.e. it is not possible to choose a specific direction for that edge without further assumptions.
@@
Finally, even though the general formulation of the discovery problem is focused on the causal graph only, causal discovery algorithms are usually designed to find a solution w.r.t. a specific set of functions [12, 83, 91, 94], e.g. non-linear equations.
@@ 2.4 Acyclicity and Faithfulness
A graphical model is said to satisfy the Markov property if the associated joint probability distribution $P(\textbf{V})$ can be decomposed recursively as:
@@
Essentially, it is assumed that probability independence ($\perp\!\!\!\perp_{P}$) implies graphical independence ($\perp\!\!\!\perp_{G}$), as stated in Equation 2.3.
@@
In order to test whether a variable $X$ is conditionally independent from $Y$ given a set Z in any probability distribution $P$ faithful to $G$, one can use the d-separation criterion which is based on the concept of blocked path.
@@
Let $G$ be a DG and $\pi$ be a path on $G$. Then, given three vertices $X$, $Y$ and $Z$ in $\pi$, we have the following:
@@
Let $G$ be a DG, $\pi$ be a path on $G$ and Z a subset of V. The path $\pi$ is blocked [28] by Z if and only if $\pi$ contains:
  • a fork $A\leftarrow B\rightarrow C$ or a chain $A\rightarrow B\rightarrow C$ such that the middle vertex $B$ is in Z, or
  • a collider $A\rightarrow B\leftarrow C$ such that middle vertex $B$, or any descendant of it, is not in Z.
The set Z d-separates $X$ from $Y$ if it blocks every path between $X$ and $Y$ [^1] In a more general setting, d-separation can be extended to set of vertices rather than just singletons [70].
[…]
Figure 2.2: In this figure, $A$ and $B$ are d-separated even without conditioning on $C$, since they form a collider. The same does not hold for $A$ and $D$, given that they form a chain by means of $C$, and therefore conditioning (i.e. setting its value) on the middle vertex $C$ d-separates them.
@@ 2.5 Equivalence Classes
The naïve representation of a DAG does not allow to convey the (lack of) knowledge that typically arise during a discovery procedure. Here, we define formally other graphical representations, along with their interpretations.
@@
The graph $G$ is a partially-directed acyclic graph (PDAG) if it can contain both undirected ()(-)( - ) and directed ()($\rightarrow$)( → ) edges.
This alternative representation allows to distinguish a cause-effect pair $(X\rightarrow Y)$ from a yet unknown relationship $(X-Y)$, where there is still uncertainty about the direction of the edge. PDAGs are also called patterns [99].
@@
In the context of PDAGs, v-structures encode the conditional independencies that shape the associated probability distribution. Any edge that would change, by either adding or removing, any v-structure when reversed is said to be a compelled edge, as in Figure 2.3. Any compelled edge, along with the underlying skeleton, is a constraint for the set of observational distributions compatible with the given PDAG. Any non-compelled edge is called reversible.
[…]
Figure 2.3: A DAG on the left and its CPDAG on the right. As we can see, both graphs have the same underlying structure (i.e. skeleton), but differ from the orientation of some of the edges. Specifically, the edges connecting $A$ to $B$ and $C$ can be rearranged to form different chains or a fork. This is not true for the others edges in the CPDAG, since they are compelled. In fact, modifying the orientation of one of them would either remove the v-structure formed by $B\rightarrow D\leftarrow C$ or introduce a new one.
@@
A PDAG $G$ is said to be completed [99] if any directed edge is compelled and any undirected edge is reversible w.r.t. MEC $[G]$.
The usual representation of a MEC is a complete partially-directed acyclic graph (CPDAG), also called essential graphs [5] or maximally oriented graphs [58]. Although the discovery problem is focused on recovering the true graph $G^{*}$ from a data set D, it is not always possible to retrieve a specific instance, but rather its MEC $[G^{*}]$.
@@
[…]
@@
The graph $G$ is a mixed graph (MG) [79, 113] if it can contain undirected ()(-)( - ), directed ()($\rightarrow$)( → ) and bidirected ()($\leftrightarrow$)( ↔ ) edges.
In mixed graphs the focus is on the edge endpoints, also called marks, rather than on the edge itself. For example, the directed edge $X\rightarrow Y$ is decomposed in two marks: the one insisting on $X\!\!\mathrel{{-}{\cdotp}{\cdotp}{\cdotp}}$ and the one insisting on $\mathrel{{\cdotp}{\cdotp}{\cdotp}{\rightarrow}}\!\!Y$. For this reason, we refer to the former as the tail mark ()(-)( - ) and the latter as the arrowhead mark (>)(>)( > ). Therefore, a bidirected edge is an edge with both marks set to arrowheads.
In a bidirected edge $X\leftrightarrow Y$, $X$ is a spouse of $Y$ and vice versa. Therefore, the set of vertices connected with a bidirected edge to $X$ is the spouse set $Sp(X)$. The graphical relationships inherited from partially directed graphs remain the same.
The fork, chain and collider patterns must be revised in the context of bidirectional edges. Let $G$ be a MG and $\pi$ a path on $G$. The pattern $X*\!\!\rightarrow Y\leftarrow\!\!*Z$ is a collider on $Y$, where ‘***’ stands for a generic mark. Any other pattern is a non-collider.
@@
The graph $G$ is a partial ancestral graph (PAG) if it can contain any combination of the following edge marks: tail ()(-)( - ), arrowhead ()($\rightarrow$)( → ) and circle ()($\circ$)( ∘ ). Moreover, let $[G]$ be the MEC associated to $G$, then:
@@
As direct consequence of this PAG definition, any circle mark present in $G$ represents a variant mark in $[G]$, as for reversible edges of CPDAGs. Thus, PAGs are the most informative representation of MECs for MAGs, hence, they satisfy the same completed definition of CPDAG.
@@
    $(X\rightarrow Y)$ $X$ causes $Y$ and $Y$ does not causes $X$, there may be an unobserved confounder,
@@
    $(X\leftrightarrow Y)$ Neither $X$ causes $Y$ nor $Y$ causes $X$, there is an unobserved confounder that causes both $X$ and $Y$,
@@
    $(X\circ\!\!\!\rightarrow\!Y)$ Either $X$ causes $Y$, or there is an unobserved confounder that causes $X$ and $Y$. While $Y$ does not causes $X$.
@@
    $(X\circ\!\!-\!\!\circ Y)$ Exactly one of the following holds: $X$ causes $Y$ or vice versa; there is an unobserved confounder that causes $X$ and $Y$; or both (1) and (3); or both (2) and (3).
Understanding which causal statement is implied by each equivalence class is fundamental for a coherent interpretation of its graphical representation.
[…]
Figure 2.4: A mixed graph on the left and one of its’ possible PAGs on the right.

### v2

@@ Abstract
Understanding the laws that govern a phenomenon is the core of scientific progress. This is especially true when the goal is to model the interplay between different aspects in a causal fashion. Indeed, causal inference itself is specifically designed to quantify the underlying relationships that connect a cause to its effect. Causal discovery is a branch of the broader field of causality in which causal graphs are recovered from data (whenever possible), enabling the identification and estimation of causal effects. In this paper, we explore recent advancements in causal discovery in a unified manner, provide a consistent overview of existing algorithms developed under different settings, report useful tools and data, present real-world applications to understand why and how these methods can be fruitfully exploited.
keywords:
causality , causal models , causal discovery , structural learning.
journal: International Journal of Approximate Reasoning
@@ 1.1 A General Overview
One of the mantras that is repeated in every statistical course is that correlation does not imply causation. This is also observed in several disciplines, such as economics [1], biology [2], computer science [3, 4] and philosophy [5]. Following [6], the main goal of a research study is often to assess the effect, if any, of an action on some outcome and not measuring a mere correlation. For example, this is true when it comes to decision making, since deciding which intervention must be taken is not straightforward and must be addressed properly to avoid any potential side effects. In order to identify and quantify a causal effect, the set of tools provided by causal discovery should be used accordingly. Here, the final goal is to decompose the total effect of an action into the causal and non-causal effects.
Causal inference, i.e. the task of quantifying the impact of a cause on its effect, relies heavily on a formal description on the interactions between the observed variables, i.e. a casual graph. Such graphical representation is naïve in its concept, yet so effective when it comes to explainability. Following [7], it boils down to connect a cause to an effect (outcome) by drawing arrows from the former to the latter, to obtain a qualitative description of the system under study. This is in stark contrast with black-box techniques, where predictions about an outcome are made with a pure data-driven approach. Indeed, these methods fall short both in terms of explainability and decision making, as stated in [8, 9, 6]. Therefore, when causality is empowered through the instrument of graphical models, it is possible to overcome the current limitations of machine learning and deep learning tools, enabling the researcher to reach a higher level of understanding.
@@
This survey is structured as follows. In Section 1, the reader is provided with a general introduction to the causal discovery problem, along with an overview of previous works on the same topic. Section 2 is devoted to provide concepts, definitions and problems that are common across different approaches presented in the following pages. Section 3 explores the first set of algorithms in the observational setting, while Section 4 relaxes the acyclicity assumption. In Section 5, the scope is extended to cover the experimental scenario, where multiple interactions with the system of interest are taken into account. Sections 6 and 7 report respectively on evaluation techniques and on practical applications of the discussed methodologies. Finally, Section 8 draws conclusions about the current landscape of causal discovery.
@@ 1.2 Related Works
To the best of our knowledge, six different surveys on causal discovery were published from 2019 to 2022. In particular, [10] acted as a meta-survey by checking the contents covered by the others concerning five topics, namely: theory, data, software, metrics and examples. A modified version of this checklist can be found in Table 1, which was adapted for a direct comparison with the structure of our survey.
@@ A Survey of Learning Causality with Data: Problems and Methods [11]
Causal Discovery Algorithms: A Practical Guide [13]
D’ya like DAGs? A survey on structure learning and causal discovery [14]
@@
While every contribution provided adequate background knowledge and theoretical definitions involving the fundamental aspects of causal discovery, only a few examples [5, 12, 15] reported evaluation datasets or metrics, and just two of them listed both [12, 10]. The landscape is even more fragmented when observed from a practical point of view: only two contributions [15, 10] presented and discussed the availability of software tools to perform the described procedures, thus hindering the applicability of causal discovery to researchers approaching this topic for the first time.
In particular, the contributions from [11, 14] provide insights on the discovery procedure using machine learning, deep learning and reinforcement learning approaches. Authors in [12] tackled the problem of recovering the causal graph from time-series datasets, while [5] restricted its attention to the most famous techniques. Moreover, [15] presented a general survey on the topic without a proper interventional section, as for [16] in the latent case. Researchers in [13] focused on a high-level overview of the methods to provide a general guide for practical applications. Finally, if the reader is interested in gaining a high-level perspective that investigates the interplay between causal inference and causal discovery, the content of [10] is preferred, opposed to our in-depth approach exclusively focused on causal discovery.
This survey is designed in the light of the above considerations and aims to guide the inexpert reader through the forest of causal graphs to avoid common pitfalls when comparing and assessing the quality of results obtained by different causal discovery algorithms. It is worthwhile to mention that this survey is different from those published from 2019 to 2022 with respect to both theory and practice. Indeed, existing surveys introduce theoretical aspects of causal discovery while only a few go into additional details. Another lack of existing surveys in term of theory is that very few of them discuss the difference between observational and interventional data. This survey has also many differences in terms of practice: i) we provide a descritpion on evaluation datasets and metrics, ii) we discuss how to tune strategies for choosing values of algorithm’s hyperparameters, iii) we report on software packages, and iv) we discuss practical applications of causal discovery methods.
@@ 2 Definitions and Notation
This section gives the main definitions, concepts and assumptions on causality, together with the associated notation. In particular, we give the definition of causal model together with the definition of causal discovery problem.
@@ 2.1 Notation
A graph $G=(\textbf{V},\textbf{E})$ is a mathematical object represented by a tuple of two sets: a finite set of vertices V and a finite set of edges $\textbf{E}\subseteq\textbf{V}\times\textbf{V}$. If not specified otherwise, this graph is intended as an undirected graph, where the undirected edge $(X,Y)$ is identical to the edge $(Y,X)$ and its graphical representation is XYX-Y.
@@
In particular, a directed edge $(X,Y)$ is graphically represented by an arrow as $X\rightarrow Y$, and induces a set of relationships between the vertices of the graph $G$. Given a vertex $X$, we denote by $Pa(X)$ its parents, i.e. the set of vertices that have an arrow into $X$, while we denote by $Ch(X)$ its children, i.e. the set of vertices that have an arrow out of $X$. Recursively, any parent and parent of a parent (child and child of a child) of $X$ is an ancestor $An(X)$ (descendant $De(X)$) of $X$.
The vertices connected to $X$ are said to be adjacent to $X$ and denoted by $Adj(X)$, while the vertices connected with an undirected edge are the neighbors $Ne(X)$. These two sets of vertices are identical in undirected graphs, but may be different in graphs with other mixed orientations.
@@
A path $\pi=(X-\cdots-Y)$ is a tuple of non repeating vertices, where each vertex is connected to the next in the sequence with an undirected edge.
@@ 2.2 Causal Model
A causal graph $G=(\textbf{V},\textbf{E})$ [8] is a graphical description of a system in terms of cause-effect relationships, i.e. the causal mechanism.
For instance, the key difference between a Bayesian network (BN) [17] and a causal Bayesian network (CBN) is the semantic interpretation of the edges. In particular, in a CBN a directed edge $X\rightarrow Y$ establishes a cause-effect relationship between the cause $X$ and its effect $Y$. Hence, reversing an edge in a BN that is not defined by a causal graph might result in the same underlying probability distribution for the pair of variables $X$ and $Y$ due to the Bayes theorem, while reversing an edge in a CBN where the graph is causal means changing the interpretation of the data generating mechanism, i.e. the causal mechanism. This concept is formalized by the following definitions.
@@
This definition is formally enforced by the causal edge assumption [9], as follows:
@@
The value assigned to each $X_{i}$ is completely determined by the function $f_{i}$ given its parents:
$X_{i}:=f_{i}(Pa(X_{i}))\qquad\forall X_{i}\in\textbf{V}.$ (2.1)
As natural consequence of Definition 2.9, we can define models that entail both the structural representation and the set of functions, i.e. the underlying causal mechanism.
@@
  • U is a set of exogenous variables, i.e. unobservable [^1] Authors in [8] define variables as exogenous if they are determined by factors outside the model, for which we choose not to explain the causes, without the precise connotation of unobservable, although typically they are. variables, where $\textbf{V}\cap\textbf{U}=\varnothing$,
@@
$M$ $=(\textbf{V},\textbf{U},\textbf{F},P)$
V $=\{X,Y,Z\}$
U $=\{U_{XY},U_{Z}\}$
F $=\begin{cases}f_{X}:X:=2U_{XY},\\ f_{Y}:Y:=X+U_{XY},\\ f_{Z}:Z:=3Y+U_{Z}\end{cases}$
$P$ $=\begin{cases}U_{XY}\sim\mathcal{N}(0,1),\\ U_{Z}\sim\mathcal{N}(0,1)\end{cases}$
Figure 2.1: The causal graph $G$ (a) of the related SCM $M$ (b). In (a) $X$ is a direct cause of $Y$ and an indirect cause of $Z$, while $Y$ is an effect, and in particular a direct effect, of $X$. An example of associated SCM is reported in (b), where the functional set F follows the causal edge assumption.
@@ 2.3 The Causal Discovery Problem
Formally, let G be the set of graphs defined over the variables V of a dataset D and $G^{*}\in\textbf{G}$ be the true but unknown graph from which D has been generated.
@@
A causal discovery algorithm is said to solve the causal discovery problem if and only if it converges to the true graph $G^{*}$ in the limit of the sample size of the datset D.
@@
A causal discovery algorithm is sound if it is able to solve the causal discovery problem, and it is complete if it outputs the most informative causal graph $G$ that can be recovered from the input dataset D, without making further assumptions.
While the criterion by which a graph is said to be “most informative” depends on the context, in general it refers to the case in which neither the presence of an edge nor its orientation can be modified further without proving more information, i.e. prior knowledge or making additional assumptions.
@@
In the following pages we will see that some algorithms are able to identify the causal graph up-to its equivalence class (Definition 2.21), meaning that setting the direction of any of the remaining undirected edges would not induce a different probability distribution, i.e. it is not possible to choose a specific direction for that edge without further assumptions.
@@
Finally, even though the general formulation of the causal discovery problem is focused on the causal graph only, causal discovery algorithms are usually designed to find a solution w.r.t. a specific set of functions [20, 21, 22, 23], e.g. non-linear equations.
@@ 2.4 Acyclicity and Faithfulness
Definition 2.15 (Markov Property).
A graph $G=(\textbf{V},\textbf{E})$ is said to satisfy the Markov property if the associated joint probability distribution $P(\textbf{V})$ can be decomposed recursively as:
@@
[…]
Essentially, it is assumed that the probability independence ($\perp\!\!\!\perp_{P}$) implies the graphical independence ($\perp\!\!\!\perp_{G}$), as stated in Equation 2.3.
@@
In order to test whether a variable $X$ is conditionally independent from $Y$ given a set Z in any probability distribution $P$ faithful to $G$, one can use the d-separation criterion, which is based on the concept of blocked path.
@@
Let $G=(\textbf{V},\textbf{E})$ be a DG and $\pi$ be a path on $G$. Then, given three vertices $X$, $Y$ and $Z$ in $\pi$, we have the following:
@@
Let $G=(\textbf{V},\textbf{E})$ be a DG, $\pi$ be a path on $G$ and Z a subset of V. The path $\pi$ is blocked [9] by Z if and only if $\pi$ contains:
  • a fork $X\leftarrow Y\rightarrow Z$ or a chain $X\rightarrow Y\rightarrow Z$ such that the middle vertex $Y$ is in Z, or
  • a collider $X\rightarrow Y\leftarrow Z$ such that middle vertex $Y$, or any descendant of it, is not in Z.
The set Z d-separates $X$ from $Y$ if it blocks every path between $X$ and YY [^2] In a more general setting, d-separation can be extended to set of vertices rather than just singletons [26]..
XXYYZZSSTTWW
Figure 2.2: In this figure, $X$ and $Y$ are d-separated without conditioning on $Z$, since they form a collider. The same does not hold for $X$ and $S$, given that they form a chain by means of $Z$, and therefore conditioning (i.e. setting its value) on the middle vertex $Z$ d-separates $X$ from $S$.
@@ 2.5 Equivalence Classes
The naïve representation of a DAG does not allow to convey the (lack of) knowledge that typically arises during a discovery procedure. Here, we define formally other graphical representations, along with their interpretations.
@@
The graph $G$ is a partially-directed acyclic graph (PDAG) if it can contain both undirected ()(-) and directed $(\rightarrow)$ edges.
This alternative representation allows to distinguish a cause-effect pair $(X\rightarrow Y)$ from a yet unknown relationship (XY)(X-Y), where there is still uncertainty about the direction of the edge. PDAGs are also called patterns [7].
@@
In the context of PDAGs, v-structures encode the conditional independencies that shape the associated probability distribution. Any edge that, when reversed, would either add or remove a v-structure is said to be a compelled edge, as in Figure 2.3. Any compelled edge, along with the underlying skeleton, is a constraint for the set of observational distributions consistent with the given PDAG. Any non-compelled edge is called reversible.
AABBCCDDEE
AABBCCDDEE
Figure 2.3: A DAG on the left and its CPDAG (Definition 2.23) on the right. As we can see, both graphs have the same underlying structure (i.e. skeleton), but differ from the orientation of some of the edges. Specifically, the edges connecting AA to $B$ and $C$ can be rearranged to form different chains or a fork. This is not true for the others edges in the CPDAG, since they are compelled. In fact, modifying the orientation of one of them would either remove the v-structure formed by $B\rightarrow D\leftarrow C$ or introduce a new one.
@@
A PDAG $G$ is said to be completed [7] if any directed edge is compelled and any undirected edge is reversible w.r.t. its MEC $[G]$.
The usual representation of a MEC is a complete partially-directed acyclic graph (CPDAG), also called essential graph [30] or maximally oriented graph [31]. Although the discovery problem is focused on recovering the true graph $G^{*}$ from a dataset D, it is not always possible to retrieve a specific instance, but rather its MEC $G^{*}]$.
@@
That is, there are no unobserved variables U that affect the behaviour of the causal mechanism generating the dataset D. If at least one latent cause exists, then V is causally insufficient, which means that there exists a non-empty set of unobserved variables U that contains at least a cause of V. In this case, $G$ is only a sub-graph of the augmented graph $G^{a}$ [24, 33] defined over $\textbf{V}\cup\textbf{U}$, as depicted in Figure 2.1. However, listing which variables are going to be included in U is not an easy task. Unless prior knowledge is available, it is generally assumed [9] that, for each variable $V_{i}\in\textbf{V}$, there exists one and only one variable $U_{i}\in\textbf{U}$ that is parent of $V_{i}$, i.e. there is an edge $U_{i}\rightarrow V_{i}$ in $G$. Hence, each endogenous variable will be influenced by one exogenous variable in the simplest scenario, while in complex settings an observed variable may share one or more unobserved parents with others variables in V, i.e. what are usually called common latent causes. On the algorithmic side, causal discovery methods, such as the Fast Causal Inference (FCI) [34], limit the search space by testing for potential unobserved variables that are common parents of two observed variables, as graphically explained in the figures in the following pages.
@@
The graph $G$ is a mixed graph (MG) [35, 34] if it contains undirected ()(-), directed $(\rightarrow)$ and bidirected $(\leftrightarrow)$ edges.
In mixed graphs the focus is on the edge endpoints, also called marks, rather than on the edge itself. For example, the directed edge $X\rightarrow Y$ is decomposed in two marks: the one insisting on $X\!\!\mathrel{{-}{\cdotp}{\cdotp}{\cdotp}}$ and the one insisting on $\mathrel{{\cdotp}{\cdotp}{\cdotp}{\rightarrow}}\!\!Y$. For this reason, we refer to the former as the tail mark ()(-) and the latter as the arrowhead mark $(>)$. Therefore, a bidirected edge is an edge with both marks set to arrowheads.
In a bidirected edge $\leftrightarrow Y, XX$ is a spouse of $Y$ and vice versa. Therefore, the set of vertices connected with a bidirected edge to $X$ is the spouse set $Sp(X)$. The graphical relationships inherited from partially directed graphs remain the same.
The fork, chain and collider patterns must be revised in the context of bidirectional edges. Let $G$ be a MG and $\pi$ a path on $G$. The pattern $X*\!\!\rightarrow Y\leftarrow\!\!*Z$ is a collider on $Y$, where ‘*’ stands for a generic mark. Any other pattern is a non-collider.
@@
The graph $G$ is a partial ancestral graph (PAG) if it contains any combination of the following edge marks: tail ()(-), arrowhead $(\rightarrow)$ and circle $(\circ)$. Moreover, let $[G]$ be the MEC associated to $G$, then:
@@
As a direct consequence of this PAG definition, any circle mark present in $G$ represents a variant mark in $[G]$, as for reversible edges of CPDAGs. Thus, PAGs are the most informative (Definition 2.12) representation of MECs for MAGs, hence, they satisfy the same completed definition of CPDAG.
@@
    $(X\rightarrow Y)$ : $X$ causes $Y$ and $Y$ does not causes $X$, there may be an unobserved confounder,
@@
    $(X\leftrightarrow Y)$ : Neither $X$ causes $Y$ nor $Y$ causes $X$, there is an unobserved confounder that causes both $X$ and $Y$,
@@
    $(X\circ\!\!\!\rightarrow\!Y)$ : Either $X$ causes $Y$, or there is an unobserved confounder that causes $X$ and $Y$. In this case, $Y$ does not causes $X$.
@@
    $(X\circ\!\!-\!\!\circ Y)$ : Exactly one of the following holds: $X$ causes $Y$ or vice versa; there is an unobserved confounder that causes $X$ and $Y$; or both (1) and (3) hold; or both (2) and (3) hold.
Understanding which causal statement is implied by each equivalence class is fundamental for a coherent interpretation of its graphical representation, as made explicit in Fig. 2.4.
AABBCCDDEE
AABBCCDDEE
Figure 2.4: A mixed graph on the left and one of its possible PAGs on the right.